Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability of hats

Expert replies
by Stockmoose16 » Wed Sep 17, 2008 11:04 am
A drawer holds 4 red hats and 4 blue hats. What is the probability of getting exactly three red hats or exactly three blue hats when taking out 4 hats randomly out of the drawer and immediately returning every hat to the drawer before taking out the next?

(a) 1/8
(b) ¼
(c) ½
(d) 3/8
(e) 7/12

My method #1: probability:

# of ways of getting 3 blue hats (w/replacement): 1/2*1/2*1/2*1/2 (the last 1/2 designates the chances of getting a red hat for the fourth selection.) = 1/16

# of ways of getting 3 red hats (w/replacement): 1/2*1/2*1/2*1/2 (the last one-half designates the chances of getting a red hat for the fourth selection.) = 1/16

My answer: 1/8


My Method #2: Combinations

4C1*4C1*4C1*4C1 = 128 (ways to select 3 blue hats)
4C1*4C1*4C1*4C1 = 128 (ways to select 3 red hats)

8C1*8C1*8C1*8C1=4096 (number of ways to pick any 4 hats)

Answer: 128/4096+128/4096
=1/8
Why are they both wrong? The OA is 1/2
Join the discussion
Source: — Problem Solving |

by mals24 » Wed Sep 17, 2008 12:47 pm
Probability of taking out 4 red hats = (1/2*1/2*1/2*1/2)*4=1/4
Probability of taking out 4 red hats = (1/2*1/2*1/2*1/2)*4 = 1/4

Total probability = 1/4+1/4 = 1/2

We multiply each probability with 4 because for each case there are 4 ways to arrange a combination of 4 hats

For the first one: BBBR, BBRB, BRBB, RBBB
For the 2nd one: RRRB, RRBR, RBRR, BRRR.
Join the discussion

by pseudononymous » Wed Sep 17, 2008 5:47 pm
Yeah, you forget to multiply by 4C3 which is 4.

4C3 * (1/2)^4 * 2 = 1/4

This is a essentially a coin-flip problem.
Join the discussion